Tuesday, April 2, 2013

Endeavor Update 13-04-02

From Chief Scientist Ray Schmitt,

Cruise Narrative, EN-522

EN-522 got underway at about 10:30 on Friday March 15 from the URI pier in Narragansett.  Seas were calm and winds were light at first, but picked up as the day wore on and we got further offshore.  Overnight we had winds in excess of 30 knots and the ship slowed and altered course to cope with the seas.  Conditions had improved by Saturday evening and we were able to make 10 knots again as we got offshore and entered the Gulf Stream.

The Thermosalinographs were running and a sampling schedule was set up for calibration bottle samples.  An additional TSG was set up and fed with a special pump and debubbling system utilizing a long hose dragging across the water.  The hose was held about 20' off the starboard side of the ship well forward of the CTD deployment area.  When the ship is going faster than a few knots the hose is sampling the upper 10 cm of the water. This system allowed us to sample any near-surface changes in salinity due to diurnal warming or rainfall.  It was set up by Julian Schanze and tuned during the transit.

After eight and a half days of steaming we arrived at the SPURS site under much improved conditions.  Winds were light and the sun was mostly bright for much of our time on site.

Sunday 3/24
The first order of business was to reset the WHOI mooring.  The releases were fired and the balls and line retrieved Sunday morning, 3/24.  We did not have to recover the whole mooring, just the synthetic line at depth.  There was a concern for chaffing at the eyesplice termination that caused us to do this line change, and indeed we did observe some chaffing.  The whole mooring operation proceeded without incident under fine weather conditions.  Completing the mooring operation allowed us to clear the decks somewhat and move the small boat to the main deck for easier deployment.

Monday 3/25
We used the small boat to launch two micro-gliders.  Seagliders were deployed from the ship.  Two WaveGliders were recovered from their holding sites near the WHOI mooring, They were snagged from the small boat then lifted aboard with the knuckle crane.  We then started steaming south to retrieve a Seaglider, the Mixed Layer Float and the drifting NOAA mooring (Pico- North).

Tuesday 3/26
This was a day of securing lost assets.  Seaglider 189 had run low on battery power and had been parked at the surface for some time.  The Mixed Layer Float deployed in September had some ballasting issues but fortunately had not drifted too far afield.  The NOAA Pico ?North mooring had broken free about a month earlier and we hoped to collect it and redeploy if possible.  All three of these were found and retrieved on Tuesday.  The small boat was used each time, in the case of the MLF and the NOAA mooring, it was to secure a means of lifting the gear aboard the ship or bring the gear to the ship, SG189 was retrieved in the small boat. That night we did a south to north U/W CTD section back toward the mooring array.

Wednesday 3/27
We performed microstructure profiling with the VMP near to ASIP and the gliders.  Several of us went over the Sarmiento for "International Salinity Summit Talks" and dinner while the Endeavor chased down the remaining WaveGlider.  This WaveGlider was steaming its programmed mission but had lost communications, so could not be commanded to its holding site.  However, it was easy to find by following its assigned track.  Endeavor returned to Sarmiento before sunset to retrieve the three scientists via small boat.  Overnight we did CTDs at the corners of the old mooring ?control volume? and UnderWay CTDs between.

Thursday 3/28
At 0800 GMT we fired the release for the NOAA Pico-North mooring to recover the release, glass balls and line in order to be able to redeploy it.  It took three hours for the balls to surface, and the initial (buoyant) line was reeled in easily.  However the polyester (heavy) line came up all snarled in large tangles.  It had clearly been sitting on the bottom, some of it was muddy.  We worked hard to untangle it but found that we had to make numerous cuts in the line to get it all aboard.  It looks like we would have to re-evaluate the plan to redeploy the mooring, though Jeff may be able to untangle and splice the line back to a usable length.  This recovery took much of the day.  That evening we deployed a Seaglider and did a CTD calibration cast near it when it dove.  We then started a CTD section along 38 W to the south with 1000m stations every 5 nm.   The Sarmiento had found a patch of fresh water intruding from the south in its SeaSoar surveys and we were planning ASIP, Glider and VMP work in a suitable fresh intrusion site so this section served in the search for the fresh intrusion.

Friday 3/29
After completing the CTD section we returned to the northwest to retrieve Seagliders 190 and 191 with the small boat.  We then steamed toward ASIP for VMP profiling.  By this time the  micro-gliders were not close to ASIP.  We worked on improving VMP profiling by adjusting winch settings and boat speed.  VMP?s continued till midnight.

Saturday 3/30
We continued the CTD section along 38 W further to the south overnight.  We then steamed to the NOAA Pico-East mooring, launched a small boat to secure a lifting tether and hoisted the buoy up at the stern A-frame using the mooring winch.  The old Prawler unit was swapped out for a new unit.  Our NOAA colleagues confirmed that it worked later in the day.  We then steamed to the WHOI Mooring and Jeff Lord visited the buoy to check meteorological equipment, one system on the buoy seems to be down.  We will likely revisit the Buoy later in the cruise once the engineers have diagnosed the problems. We then steamed to recover the two microgliders, Helo and Saul, securing them just before sunset.  Plans were made to re-deploy with ASIP on Sunday at a site to the south-south west so a north to south CTD section was done overnight along 38 25' W.

Sunday 3/31/
We stopped the section at 23 50' N and steamed a few miles east to retrieve one of last year?s ARGO floats that had developed a small leak.  We then steamed to the ASIP position and deployed Helo and Saul quite near it.  We then deployed the VMP first steaming north away from the cluster of instrument then approaching again from the north along a parallel section about 1 km west of their position.  This was continued till 2100 (local) then we restarted the CTD section further south along 38 25'W.

Monday 4/1
The CTD section along 38 25'W was continued, reaching 23 10'N.  We then steamed north back to the ASIP/Glider site for further VMP profiling.  After several hours the Sarmiento steamed into site and the chief scientists discussed plans over VHF.  The Gliders were recovered at about 16:30, while Sarmiento recovered ASIP.  We then began steaming to the north for a CTD section north along 38 W starting near the NOAA Pico-North mooring site.  The NOAA mooring redeploy is planned for tomorrow morning.

No comments:

Post a Comment